LCD to determine which fraction is greater.
5

12
, 7

11


Answer by stanbon(75887)

(5/12) R (7/11)
Cross-multiply
55 R 84
----
R is "less than"
So (5/l2) is less than (7/11)
-----------------------------------
Using the lcd:
The lcd is 12*11
Rewrite each fraction the the lcd as its denominator:
(5/12) = (5*11/12*11) = 55/lcd
---------------------
(7/11) = (7*12/11*12) = 84/lcd
------------------------------------
55 is less than 84
So, (5/12) is less than (7/11)
==================================
Cheers,
Stan H.